OpenStudy (anonymous):

Provide an example of an arithmetic series which totals zero. Using complete sentences, explain how you created the example.

OpenStudy (amistre64):

infinite series or finite?

OpenStudy (amistre64):

a + d(1) a + d(2) a + d(3) a + d(4) ------- 4a + d(10) = 0 d(10) = -4a d = -2a/5 say a =, i dunno, 2 for example 2 - 4/5 = 6/5 2 - 8/5 = 2/5 2 -12/5 = -2/5 2 -16/5 = -6/5 ------ 0

OpenStudy (anonymous):

oh, how did you create that example?

OpenStudy (anonymous):

a series example

OpenStudy (amistre64):

just start in the middle and add and subtract yor way out

OpenStudy (anonymous):

Which one is the series?

OpenStudy (cwrw238):

a simple way is to choose 4 numbers each side of 0 which have a common ratio such as -2, -1, 0, 1, 2

OpenStudy (amistre64):

......a-d(n),..., a ,..., a + d(n),...

OpenStudy (amistre64):

yeah, but a common difference instead :)

OpenStudy (cwrw238):

* common difference

OpenStudy (cwrw238):

right amistre!!

OpenStudy (amistre64):

the zero can be excluded, or kept intact ...

OpenStudy (amistre64):

my original took the general equation to define a means to determine a suitable series; but as it turns out .... it just amounts to negations

OpenStudy (cwrw238):

choose a negative first term and a common difference say -6 and 4 then add terms so as to make the sum zero- -6, -2, 2, 6

OpenStudy (anonymous):

Thanks guys